32g of CaC 2 react at 27°C with excess of water to produce acetylene gas. What is work done? a)-2400 Cal b)-300 Cal c)-300 J d)-2400 L-atm Please explain with detailed explanation

32g of CaC2 react at 27°C with excess of water to produce acetylene gas. What is work done?


a)-2400 Cal           b)-300 Cal           c)-300 J            d)-2400 L-atm


Please explain with detailed explanation

Grade:upto college level

1 Answers

Saurabh Gupta
10 years ago

Write down the balanced reaction.

 

CaC2 + 2 H2O (excess) → C2H2 + Ca(OH)2

Molar Mass of CaC2 = 40 + (12*2) = 40 + 24 = 64 g/mol

No. of Mole of CaC2 = 32/64 = 0.5 mole

 

Now, Work Done = ΔngRT,

where, Δng is difference in no. of (gaseous) mole of reactant and product;

R is the Universal Gas Constant;

T is the absolute temperature.

 

Thus, W = (-0.5)(8.314)(300) = 1247.1 J

or W = (-0.5)(0.0821)(300) = 12.315 L-atm

or W = (-0.5)(1.987)(300) = 298.05 cal ≈ 300 cal

 

Hence option (b) is correct.